### v1

# From Word to World: Evaluate and Mitigate Culture Bias via Word Association Test

Abstract
The human-centered word association test (WAT) serves as a cognitive proxy, revealing sociocultural variations through lexical-semantic patterns. We extend this test into an LLM-adaptive, free-relation task to assess the alignment of large language models (LLMs) with cross-cultural cognition. To mitigate the culture preference, we propose CultureSteer, an innovative approach that integrates a culture-aware steering mechanism to guide semantic representations toward culturally specific spaces. Experiments show that current LLMs exhibit significant bias toward Western cultural (notably in American) schemas at the word association level. In contrast, our model substantially improves cross-cultural alignment, surpassing prompt-based methods in capturing diverse semantic associations. Further validation on culture-sensitive downstream tasks confirms its efficacy in fostering cognitive alignment across cultures. This work contributes a novel methodological paradigm for enhancing cultural awareness in LLMs, advancing the development of more inclusive language technologies.
From Word to World:
Evaluate and Mitigate Culture Bias via Word Association Test
Xunlian Dai1,2, Li Zhou1thanks: Corresponding author, Benyou Wang1, Haizhou Li1,3 1The Chinese University of Hong Kong, Shenzhen 2Wuhan University of Technology 3Shenzhen Research Institute of Big Data, lizhou21@cuhk.edu.cn
1 Introduction
Word associations, often perceived as commonplace neural activities Pulvermüller (1999); Anderson et al. (2017); Khanna et al. (2024), are deeply rooted in lived environments and personal experiences. The neurocognitive research Schneider et al. (2024) find out that transient human associations reflect not only situational factors but also ingrained in sociocultural identities. Cross-cultural WAT also unveil how culture modulates perceptual and interactive patterns Szalay and Deese (2024).
For instance, as shown in Figure 1, when prompted with “red,” beyond general associative words such as “blue” and “color”, which reflects semantic relations, British respondents tend to activate idiomatic expressions like see red (anger), while Australians are more likely to mention bureaucratic metaphors such as red tape(bureaucracy). This cross-cultural divergence in word associations reveals rich cognitive-semantic information embedded within linguistic conditioning.
Figure 1: Cross-cultural comparison of word associations for the cue “Red” between LLM predictions and human responses.
Multilingual LLMs exhibit persistent cultural biases, ranging from stereotypical associations Abid et al. (2021); Bano et al. (2025) to value misalignments Masoud et al. (2023); Cao et al. (2023a); Jiang et al. (2024). Prior research has largely concentrated on addressing these biases by aligning multicultural cognition through prompt-based methods Choenni and Shutova (2024); Wang et al. (2024); Sato et al. (2024), which depend on explicitly providing cultural context to elicit targeted responses. On the other hand, fine-tuning approaches Xu et al. (2024); Yao et al. (2025) are knowledge-driven but still require explicit cultural settings during inference. In contrast, our approach implicitly enhances the model’s cultural awareness and sensitivity by simulating human word association tasks and learning semantic spaces of cultural preferences within the model itself, directly capturing latent cultural relationships. Our contributions can be summarized as follows:
  • We design an LLM-adaptive, free-relation word association task and a quantitative evaluation metric to assess the cross-cultural cognitive abilities of LLMs (§3).
  • We introduce CultureSteer, an innovative approach that integrates a culture-aware steering mechanism to guide semantic representations toward culturally specific spaces (§ 4).
  • Our experiments show that LLMs exhibit cultural bias in word association tasks, while our model is able to effectively mitigate this bias, outperforming prompt-based methods (§ 5).
  • Further analysis demonstrates the effectiveness of the culture-aware steering mechanism and the generalizability of the model (§ 6).
2 Related Work
Cultural Bias and Alignment in LLMs
Extensive research has investigated cultural biases in LLMs across domains such as value preference Benkler et al. (2023); Cao et al. (2023a); Jiang et al. (2024); Durmus et al. (2023), knowledge perception Palta and Rudinger (2023); Shen et al. (2024), and moral measurement Jinnai (2024); Rao et al. (2023, 2024), revealing persistent Western-centric value preferences and regionally divergent commonsense knowledge. To improve cultural perception, current approaches predominantly employ prompt engineering Wang et al. (2024); AlKhamissi et al. (2024); Zhou et al. (2025); Masoud et al. (2025), explicitly embedding cultural context in prompts to evoke culturally sensitive responses. Alternatively, fine-tuning methods Yao et al. (2025); Xu et al. (2024) have been proposed to enhance cultural alignment via few-shot learning or cross-cultural similarity modeling. Our approach stands out by proposing a novel word-level perspective through free word associations, directly integrating cultural knowledge into the model’s internal representations via associative fine-tuning, thereby achieving deeper cultural alignment beyond prompt-based methods.
Word Association in LLMs
Recent work adopts word association tasks to evaluate the capabilities of LLMs across various domains. For instance, uncovering gender stereotypes Abramski et al. (2024b) and assessing color associative perception Fukushima et al. (2024). These studies primarily focus on word association mappings under specific relationships, emphasizing explicit word generation and partially relying on manual interpretation during evaluation Abramski et al. (2024a); Vintar et al. (2024).
Steering Control of LLMs
Steering methods aim to control the internal activation states of LLMs at inference time to influence the generation style of their outputs Bo et al. (2025). These have been applied to reduce harmful responses Arditi et al. (2024); Cao et al. (2024) and perform style transfer Song et al. (2025); Konen et al. (2024). Our approach introduces a steerable layer that modulates associative outputs to reflect culture-specific cognitive patterns in its lexical associations.
3 Word Association Test Task
Word Association Test Woodworth and Wells (1911); Gough (1976) is a psychological assessment that reveals subconscious thoughts, emotional states, or cognitive patterns through an individual’s quick associative responses to stimulus words. The cultural background influences the associated words and response patterns, with different cultures often producing vastly different associations to the same stimulus, reflecting the profound impact of culture on language, thinking, and emotions.
3.1 Human-centered task
The WAT for human participants is conceptualized as an open-ended activity in a free-relation format, allowing for flexible and diverse associations that more accurately reflect human cognitive processes. [^1] https://smallworldofwords.org Before conducting the WAT, demographic information, such as age, gender, education, native language, and region, is collected for each participant. Given a cue word $w$, the participant $u$ is asked to provide the first word that comes to mind and may optionally add a second and third word sequentially, resulting in up to three associated words, represented as $u_{w}^{c}=\left\{a_{1}^{u},a_{2}^{u},a_{3}^{u}\right\}$, where $c$ denotes the participant’s cultural background. Within this cultural context, the collective set of associative words for a given cue word $w$, aggregated across all participants $U_{w}^{c}=\left\{u_{w}^{c}\right\}$, is represented as $A_{w}^{c}=\left\{\left(a_{1},f_{1}\right),\cdots,\left(a_{n},f_{n}\right)\right\}$. Here, $f_{i}\geqslant f_{i+1}$, and $f_{i}$ represents the frequency with which the associative word $a_{i}$ appears in $U_{w}^{c}$. Across different cultural contexts, the associative words in $A_{w}^{c}$ may exhibit nuanced variations in both semantic content and frequency distribution.
3.2 LLM-adaptive task
Directly employing the human-centered task paradigm to conduct WAT with LLMs proves insufficient for effectively capturing variations across entire cultural groups. Inspired by Fierro et al. (2024) and Zhou et al. (2025), we adapt this task into a word prediction framework, designed as a retrieval process for candidate associative words. Specifically, we employ a template $q$, such as “When {cue_word} is mentioned, people often associate it with the following words:”, to prompt the LLM and elicit its predictions of associated words. The input is represented as $x=q(w)$, and the output is denoted as $o=M(q(w))$, where the generalized function $M(\cdot)$ represents the inference process of the LLM.
Unlike the human-centered task, which gathers associative words directly from participants, our method computes the probability of each candidate associative word being predicted by the LLM and ranks them based on their relative probabilities. [^2] This can be viewed as an open-ended multi-choice question-answering task without predefined options. To address potential uninformative prefixes in the LLM’s next-token predictions (e.g., phrases like ‘I think’), we move beyond relying solely on the likelihood of a candidate associative word appearing as the immediate next token. Instead, we compute its probability within a fixed window of next-token predictions. By evaluating the candidate word’s probability at each position within the prediction window, we select the maximum value as its final probability, $P_{w}\left(a_{i}\right)$. The mathematical formulation of this process is as follows:
$P_{w}\left(a_{i}\right)=\overset{}{\max_{m\in\left[0,k-t\right]}}\left(\frac{1}{t}\sum_{j=0}^{t-1}{p\left(\hat{y}_{m+j}=a_{i}^{j}\right)}\right)$ (1)
where $k$ is the size of the fixed prediction window, $t$ is the number of sub-tokens in the candidate word $a_{i}$, $a_{i}^{j}$ is the $j$-th sub-token, and $\hat{y}_{m+j}$ is the token predicted at position $m+j$. Based on these probabilities, we construct the predicted ranked list of associative words, denoted as $\hat{A}_{w}=[\hat{a}_{0},\hat{a}_{1},\dots]$, where $P_{w}\left(\hat{a}_{i}\right)\geqslant P_{w}\left(\hat{a}_{i+1}\right)$.
Figure 2: The framework of CultureSteer model. Top: model pipeline; Bottom: training and inference process based on WAT.
3.3 Evaluation metric
To evaluate the alignment between human cultural associations $A_{w}^{c}$ and the LLM’s predicted associations $\hat{A}_{w}$, we propose a novel evaluation metric, Position-Weighted Recall (PWR@K), which is an extension of Top-K Recall (R@K).
Baseline Metric
Given that WAT in LLMs is a retrieval task involving a large number of truly relevant associative words and a ranking process close to full ranking, we adopt R@K as the baseline evaluation metric, emphasizing full coverage. Formally, it is defined as:
$\text{R@K}&={\frac{\sum_{i=1}^{K}{\mathbb{I}\left(\hat{a}_{i}\in A_{w}^{c}\right)}}{N}}\\ &=\frac{\sum_{i=1}^{N}{\mathbb{I}\left(a_{i}\in\text{Top-K}\left(\hat{A}_{w}^{c}\right)\right)}}{N}\end{aligned}$ (2)
where $N=\left|A_{w}^{c}\right|$ denotes the total number of truly relevant associative words, and $\mathbb{I}(\cdot)$ is the indicator function.
Proposed Metric
While R@K effectively measures coverage, it fails to account for ranking order. In our task, higher-ranked words in $A_{w}^{c}$ signify stronger associative strength and thus hold greater importance. To address this limitation, we propose PWR@K, which incorporates positional weighting by assigning higher weights to words ranked earlier, specifically using the inverse of their position as the weight. It is formally defined as:
$\text{PWR@K}=\frac{\sum_{i=1}^{N}\frac{1}{i}\cdot\mathbb{I}(a_{i}\in\text{Top-K}(\hat{A}_{w})}{\sum_{i=1}^{N}\frac{1}{i}}$ (3)
where the denominator normalizes the positional weights across all $N$ positions.
4 CultureSteer: Modeling and Controlling Cultural Awareness in LLMs
Inspired by the findings of LM-Steer Han et al. (2024), which suggest that, with certain assumptions, shifting styles in language models is equivalent to a linear transformation in the word embedding space, we extend this perspective to the cultural domain. Specifically, we hypothesize that different cultures define distinct semantic association spaces shaped by cultural preferences, which can be modeled through culture-specific linear transformations within the embedding space.
Based on this, we introduce the CultureSteer model, which leverages both human-centered and LLM-adaptive WAT tasks to (1) equip LLMs with human-like word association perception, thereby enhancing their cognitive abilities, and (2) enable the word association capabilities of LLMs to align across different cultures. The overall architecture of the model is illustrated in Figure 2.
Cultural Control Paradigms
Previous approaches primarily control cultural preferences explicitly at the input level by defining the cultural context $c$ through textual prompts. The output in these methods can be denoted as:
$o=M\left(q\left(w,c\right)\right)$ (4)
In contrast, our model performs post-processing on the LLM’s output, mapping the semantic representation space to a specific cultural perception space. This operation is represented as:
$o=\mathrm{Steer}\left(M\left(q(w)\right),c\right)$ (5)
Cultural-Aware Steering Mechanism
Building on this paradigm, we detail the CultureSteer framework, focusing on how semantic representations are steered into culturally-specific spaces. First, we obtain the semantic representation $\boldsymbol{h}=M\left(q\left(w\right)\right)$, which encodes the general meaning without any culture-specific preferences. To model culture-specific semantic associations, we apply a transformation to the base semantic representation $\boldsymbol{h}$. Specifically, the cultural preference adjustment $\bigtriangleup\boldsymbol{h}^{\left(c\right)}$ is computed as:
$\bigtriangleup\boldsymbol{h}^{\left(c\right)}=\boldsymbol{h}\cdot\left(\boldsymbol{c}\odot\mathbf{W}\right)=\boldsymbol{h}\cdot\mathbf{W}_{c}$ (6)
Here, $\boldsymbol{c}\in\mathbb{R}^{\left|C\right|}$ is the cultural control vector, where each element corresponds to a specific culture. In our work, we use a one-hot encoding for different cultural contexts. The operation $\boldsymbol{c}\odot\mathbf{W}$ selects the relevant cultural subspace from $\mathbf{W}$ called $\mathbf{W}_{c}$, which contains learnable parameters that encode culture-specific semantic directions. The final culture-aware representation is then computed as:
$\tilde{\boldsymbol{h}}^{\left(c\right)}=\boldsymbol{h}+\epsilon\bigtriangleup\boldsymbol{h}^{\left(c\right)}$ (7)
where $\epsilon$ is a scaling factor that modulates the influence of the culture-specific adjustment. [^3] We set $\epsilon=1e-3$ as the default steering value Han et al. (2024). The adjusted representation $\tilde{\boldsymbol{h}}^{\left(c\right)}$ is then used to predict the associated words for a given query, considering the specified cultural context.
Training and Inference Process
During training, we use the human-centered WAT pattern to train the CultureSteer model SFT, where the true human associations are fed back into the model at each time step $t$. The loss function, which minimizes the discrepancy between the predicted associations $\tilde{a}_{t}^{u}$ and the true human responses $a_{t}^{u}$. used is cross-entropy During the inference phase, we employ the LLM-adaptive WAT task to evaluate the LLM’s cross-cultural word association ability, using the proposed metric PWR@K.
Llama Qwen
PWR@K K=3 K=5 K=10 K=20 PWR@K K=3 K=5 K=10 K=20
Baseline 12.37 19.57 29.25 42.54 Baseline 15.22 19.05 26.18 35.90
CSP 30.30 37.61 48.95 60.95 CSP 20.34 28.19 39.05 50.65
CCT 25.92 32.98 44.05 55.54 CCT 22.59 30.51 42.31 54.76
USA CultureSteer 38.14 50.30 67.58 80.13 CultureSteer 36.34 46.06 59.78 72.05
Baseline 10.69 16.53 24.96 35.92 Baseline 12.43 16.16 22.05 29.97
CSP 24.31 30.64 39.91 49.27 CSP 16.43 22.14 30.58 39.86
CCT 21.58 27.74 35.70 44.96 CCT 18.17 24.27 33.54 42.72
UK CultureSteer 29.19 38.84 51.45 61.88 CultureSteer 28.23 34.88 46.10 57.53
Baseline 9.53 15.62 23.25 33.67 Baseline 11.87 15.07 21.35 28.41
CSP 23.17 28.62 36.78 45.88 CSP 15.58 21.21 29.02 37.60
CCT 20.21 25.64 33.37 42.41 CCT 17.16 22.67 31.61 40.63
OC CultureSteer 27.56 35.44 47.65 58.77 CultureSteer 25.05 32.22 41.55 52.22
Baseline 7.31 11.03 18.43 24.01 Baseline 8.61 11.65 17.16 24.11
CSP 10.52 15.83 23.34 29.81 CSP 15.56 21.59 31.77 42.81
CCT 9.71 14.08 21.02 28.00 CCT 12.32 17.71 26.57 36.85
CN CultureSteer 12.77 19.04 27.26 33.53 CultureSteer 21.78 30.13 44.88 58.76
Baseline 9.98 15.69 23.97 34.04 Baseline 12.03 15.48 21.69 29.60
CSP 22.08 28.18 37.28 46.48 CSP 16.98 23.28 32.61 42.73
CCT 19.36 25.11 33.54 42.73 CCT 17.56 23.79 33.51 43.74
Average CultureSteer 26.92 35.91 48.49 58.58 CultureSteer 27.85 35.82 48.08 60.14
Table 1: Overall comparison results: (1) LLMs display cultural biases in the WAT task; (2) CultureSteer enhances LLMs’ cultural awareness and achieves the best performance.
5 Experiment
5.1 Dataset Creation
Data Source
We utilized the officially released datasets from the Small World of Words Project (SWOW) [^4] Designed as a human-centered Word Association Task (WAT).. For our study, we selected data in two languages: English and Mandarin Chinese, referred to as SWOW-EN De Deyne et al. (2019) and SWOW-ZH Li et al. (2024), respectively. The SWOW-EN dataset comprises over 3 million responses collected from more than 90,000 participants across various English-speaking countries or regions, covering more than 12,000 cues. Moreover, the SWOW-ZH dataset contains over 2 million responses corresponding to 10,192 cues, with participants primarily from the Chinese mainland.
Culture Selection
To ensure sufficient participant representation across different cultures, we first filter the three most prominent cultural groups from SWOW-EN: the United States (USA), the United Kingdom (UK), and Oceania (OC) [^5] OC includes participant responses from Australia and New Zealand, which are combined due to their relatively small participant numbers, geographical proximity, and shared historical background.. Additionally, SWOW-ZH serves as the representative for China (CN). As a result, we construct datasets representing four distinct cultural groups: USA, UK, OC, and CN.
Data Preprocessing
To ensure fairness in evaluating word association perception across different cultures, we utilize the Intercontinental Dictionary Series (IDS) jung Yu and Wang (2023) to align the cue words between English and Chinese, ultimately retaining 881 cue words that are common across all four cultural groups. Given the substantial differences in participant numbers across cultures, which lead to significant variation in the number of associated words ($A_{w}^{c}$), we standardize the $A_{w}^{c}$ lists by truncating them to the size of the culture with the fewest associated words. Using the IDS framework, we categorize these cue words into 22 semantic categories based on their conceptual meanings. Examples of the final preprocessed data and statistical details are provided in Appendix B.2, Table 4 and Table 5. Finally, we split the cue words in each semantic category into 70% for training and 30% for testing.
5.2 Experimental Setup
Figure 3: Fine-grained performance comparison across 22 semantic classes in the test set with PWR@20. Red denotes Global Knowledge, green denotes Perceptual Experience, and blue denotes Cultural Ideologies. Other PWR@K (K=3, 5, 10) results are shown in Appendix D.
We conduct experiments using Llama3.1-8B Grattafiori et al. (2024) and Qwen2.5-7B Team (2024) to simulate the WAT. First, we evaluate the base versions of these models on word association tasks using a base prompt without any cultural conditioning, aiming to assess their human-like associative cognition and the default cultural representation they exhibit. Subsequently, we build CultureSteer based on these two models to enhance their cross-cultural perception capabilities. For the LLM backbone, we adopt LoRA Hu et al. (2022), a parameter-efficient tuning method, implemented via LLaMAFactory, and train the models on a single A100 GPU. Furthermore, we introduce two explicit culture-aware prompting strategies as compared methods: Culture-Specific Prompt (CSP) Zhou et al. (2025), which defines a specific cultural context in the prompt, and Cross-Culture Thinking (CCT) Xu et al. (2025), which prompts LLMs to consider the relationships between the given culture and other cultures. Detailed prompt settings are provided in Appendix B.1.
TopK Model Culture w/o $\mathbf{W}$ $\mathbf{W}_{USA}$ $\mathbf{W}_{UK}$ $\mathbf{W}_{OC}$ $\mathbf{W}_{CN}$ PWR@K Model Culture w/o $\mathbf{W}$ $\mathbf{W}_{USA}$ $\mathbf{W}_{UK}$ $\mathbf{W}_{OC}$ $\mathbf{W}_{CN}$
K=3 Llama USA 36.28 38.14 36.21 36.32 36.32 K=5 Llama USA 47.96 50.30 47.99 48.07 47.95
UK 27.80 27.78 29.19 27.96 27.84 UK 37.03 37.03 38.84 37.08 37.03
OC 26.48 26.48 26.36 27.56 26.49 OC 33.57 33.55 33.73 35.44 33.56
CN 12.23 12.23 12.23 12.21 12.77 CN 18.30 18.30 18.36 18.36 19.04
Qwen USA 37.14 38.04 34.83 34.72 34.74 Qwen USA 47.24 47.91 43.87 43.83 43.89
UK 26.45 26.72 27.84 26.79 26.77 UK 34.01 33.21 35.69 33.26 33.27
OC 26.82 25.23 25.20 25.05 25.19 OC 31.35 31.97 31.97 32.22 32.01
CN 21.49 19.54 19.34 19.35 21.78 CN 24.81 24.44 24.15 24.14 30.13
K=10 Llama USA 64.43 67.58 64.50 64.56 64.40 K=20 Llama USA 76.11 80.13 76.20 76.05 76.08
UK 48.88 49.02 51.45 48.92 48.85 UK 58.87 58.97 61.88 58.77 58.87
OC 45.09 45.10 44.96 47.65 45.05 OC 55.89 55.89 56.04 58.77 55.89
CN 26.12 26.12 26.12 26.12 27.26 CN 32.09 32.09 32.09 32.09 33.53
Qwen USA 56.19 61.88 56.81 56.83 56.86 Qwen USA 68.34 73.94 68.82 68.92 68.89
UK 43.61 43.92 46.45 44.02 43.87 UK 54.58 54.78 56.66 54.74 54.78
OC 41.19 41.28 41.14 41.55 41.11 OC 50.73 50.70 50.73 52.22 50.85
CN 33.53 33.01 32.62 32.73 44.88 CN 44.34 44.51 44.59 44.57 58.76
Table 2: Results of controlled experiments on cultural-aware steering mechanism: performance declines without cultural preference semantic learning or with mismatched cross-cultural semantic space steering, highlighting the existence and variability of cultural semantic preferences.
5.3 Overall Results
LLMs Exhibit Cultural Biases in the WAT Task
The overall comparison results are presented in Table 1, where the “Baseline” denotes the default cultural preferences exhibited by the LLMs. Both models consistently reveal a cultural preference pattern: USA > UK > OC > CN. [^6] To demonstrate that performance differences are due to cultural factors rather than the language used in prompts, we conducted an ablation study. Details are provided in Appendix C. This suggests that, in terms of word-level cognitive abilities, LLMs exhibit a stronger inclination toward American culture, aligning with findings from prior studies Cao et al. (2023b); Myung et al. (2024); Zhou et al. (2025). Overall, Llama demonstrates stronger human-like word-level cognitive abilities compared to Qwen, primarily due to its significantly better performance in English-speaking cultural contexts. In contrast, the performance gap between the two models is relatively small in CN. Interestingly, when evaluating PWR@3, Qwen outperforms Llama across all cultural contexts, suggesting that Qwen excels at predicting highly associated words. However, as K increases, Llama surpasses Qwen, indicating that Llama possesses a more comprehensive associative cognition across a broader range of related words.
CultureSteer Enhance LLM Cultural Awareness
At the methodological level, CultureSteer consistently outperforms both prompt-based methods and the base models (Llama and Qwen) across all cultural settings. This demonstrates that the model does not rely on explicit textual configurations. Instead, it leverages linear transformations to implicitly align with the semantic space of specific cultural preferences, effectively adjusting the default semantic space to enhance its cultural awareness. Interestingly, for the Llama model, the CSP method consistently outperforms the CCT method across all cultural contexts. In contrast, for the Qwen model, the CSP method performs better only in the Chinese cultural context, while the CCT method excels in the other three, particularly in English-speaking settings. This suggests that Llama is more effective with culture-specific prompts tailored to a single context, whereas Qwen demonstrates stronger cross-cultural reasoning, especially for the English-speaking environments. This demonstrates that explicitly simulating cultural contexts through prompt design highlights the sensitivity of LLM performance to specific prompt configurations. Moreover, it further validates the effectiveness of CultureSteer’s implicit learning paradigm in enhancing cultural awareness.
5.4 Fine-Grained Results
We visualize the fine-grained performance comparison across 22 semantic classes in the test set, as shown in Figure3. The results reveal notable differences among baseline models in the Cognition and Value classes, with Qwen demonstrating value orientations more aligned with Chinese culture, while Llama exhibits stronger alignment with English cultural values. Furthermore, based on their semantic meanings, these 22 classes are categorized into three groups: Global Knowledge, Perceptual Experience, and Cultural Ideologies. CultureSteer improves cultural awareness in both models with domain-dependent effectiveness. While global knowledge categories (e.g., World, Animals) remain difficult to align culturally, perceptual experience categories (e.g., Action, Space) show notable gains, especially within English-speaking cultures. Llama still exhibits gaps between English and Chinese contexts after steering, whereas Qwen achieves more balanced cross-cultural performance, nearing parity in cultural ideologies (Values, Society). Overall, CultureSteer mitigates cultural biases and enhances associative abilities but cannot fully overcome inherent differences in models’ cultural perception.
6 Analysis
6.1 Culture-Specific Semantic Learning Matters
To validate the effectiveness of our proposed cultural-aware steering mechanism, we conduct two controlled experiments: one without cultural preference semantic learning in training and one with cross-cultural semantic space steering in inference. The results are shown in Table 2. We observe a general decline in performance when the cultural semantic preference space is not learned or when a mismatched cultural semantic space is used for steering. This finding confirms the presence of cultural semantic preferences, highlights the differences across cultural semantic spaces, and underscores the importance of implicitly learning these preferences. For more detail, on LLama, no significant performance differences are observed when comparing the absence of cultural preference semantic learning with the use of mismatched cross-cultural semantic space steering. However, on Qwen, mismatched cultural semantic spaces have a negative impact on performance when evaluated at $K\leq 10$. Interestingly, as the evaluation range increases to $K=20$, a slight improvement in performance is observed, which suggests that broader evaluation scopes may capture additional relevant associations despite the mismatch, partially offsetting the negative effects.
Llama Qwen
Baseline CultureSteer Baseline CultureSteer
WVS USA 53.64 51.11 47.66 51.43
UK 51.69 53.89 48.33 51.57
OC 51.43 53.32 46.63 52.35
CN 49.11 53.11 46.87 49.01
BLEnd USA 46.29 50.47 62.73 68.91
UK 43.18 48.86 57.89 62.86
CN 41.65 47.93 52.84 56.55
Table 3: Performance Comparison on Cross-Cultural Benchmarks: WVS and BLEnD. Bold values indicate superior performance under each model. In WVS, we evaluated the Australian Q&A data as the result for OC, while BLEnD lacks data from Australia and New Zealand, so no evaluation was conducted.
6.2 Generalization on Cultural Tasks
To assess whether CultureSteer’s cultural awareness and cognitive modeling capabilities benefit downstream culture-related tasks, we perform experiments on two datasets: the World Values Survey (WVS) Xu et al. (2024), which captures global perspectives on values and beliefs, and BLEnD Myung et al. (2024), which focuses on everyday knowledge in diverse cultures. The experimental results are shown in Table 3. CultureSteer consistently outperforms across all cultural settings on both tasks, except for the USA culture in the WVS task, indicating that our method enhances LLMs’ cultural awareness through word-level interventions. The performance gains on both WVS and BLEnD confirm that mitigating cultural bias in word association tasks establishes a foundational cognitive structure for broader cultural competence. By grounding conceptual mappings in cultural contexts through word association mechanisms, CultureSteer activates latent cultural perception capabilities in language models, surpassing superficial mitigation of task-specific biases. This validates that word association analysis serves as an effective entry point for enhancing the generalization capacity of multicultural LLMs.
6.3 Case Study
To gain deeper insights into how CultureSteer differentially influences the probability distributions of $\hat{A}_{w}^{c}$, we perform a targeted case study focusing on the cue word “red” across three English-speaking cultures: USA, UK, and Oceania (OC). [^7] Chinese (CN) is excluded to avoid confounding linguistic differences across language families. We compute the probability differences between CultureSteer and the baseline LLM version for $A_{\text{red}}^{c}$, as illustrated in Figure 4. The results reveal notable probability increases for both culture-specific words (red marks) and general words shared across cultures. This indicates that the model learns not only universal word associations but also culture-specific ones, validating the effectiveness of the CultureSteer.
7 Conclusion
Our research aims to utilize the insights from WAT to evaluate and mitigate cultural biases in LLMs. We proposed a novel metrics for assessing the association abilities of LLMs through word prediction task. Additionally, we introduced CultureSteer that leverages word association patterns to align LLMs’ cultural cognition by guiding sematic representations toward culturally specific spaces and validated effectiveness of this method in downstream tasks. Although we did not explicitly prescribe the specific associative logic underlying word associations, we instead enabled the model to infer these connections through contextual prompts. Remarkably, our training methodology successfully elicited the model’s cross-cultural word association capabilities. Future research could concentrate on refining these associative patterns to further augment models’ cultural sensitivity and semantic comprehension.
Figure 4: Probability differences between CultureSteer and vanilla models across USA, UK and OC. Red marks denote culture-specific words, while gray marks represent general words common across cultures.
Limitation
Our research is constrained by its focus on a limited number of languages and cultures when examining word association differences, primarily due to the reliance on word association test datasets sourced from human participants, which poses significant challenges to scalability. Moreover, our control vector mechanism treats cultures as independent entities, without adequately considering potential relationships and interactions between them. Future work could address these limitations by developing scalable methods for collecting word association data, expanding the linguistic and cultural scope, and incorporating inter-cultural connections to provide a more comprehensive and nuanced perspective.
Appendix A Data Preparation
A.1 Few example of IDS chapters
We list the whole 22 chapters of IDS and a few example words, shown in Table 4
Chapter Examples(English)
Physical world flame, ice, light
Kinship they, girl, male
Animals bee, ant, fox
The body blind, wound, dream
Food and drink fruit, mill, dish
Clothing and grooming headband, coat, tailor
The house brick, window, yard
Agricultural and vegetation crop, plow, garden
Action and technology wash, untie, glass
Motion harbor, carriage, road
Possession market, trade, tax
Spatial relations wide, south, far
Quantity second, some, more
Time often, age, summer
Sense perception see, color, heavy
Emotion and values joyful, love, hate
Cognition know, learn, pupil
Speech and Language ask, refuse, paper
Society and politics master, friend, custom
Warfare and hunting fight, soldier, shoot
Law court, accuse, steal
Religion and belief bless, heaven, ghost
Table 4: The 22 chapters names with few sample words
A.2 Data Cleaning
We filter SWOW with IDS by the following steps:
  • For one-to-many alignments in IDS, we retain only the first word in the list of multiple terms.
  • Align the cue words between English and Chinese in IDS cross-culture word paris
  • Remove abnormal characters and meaningless responses, such as ’#Missing’ or ’?’.
  • Filter out responses containing multiple words in English and those with more than four characters in Chinese.
  • Truncate the association words for each cue word to the number of association words corresponding to the culture with fewest association words.
A.3 Data Overview
The data overview(frequency) of our dataset has been shown in Table 5. By truncating the association word, we ensure that the number of tuples for each association sorting $A_{w}$ corresponding to each cue word is consistent across cultures.
Chap Cue Asso (avg.) Users
USA UK OC CN
World 52 15.96 3895 645 513 1995
Kinship 34 14.21 2246 342 266 1282
Animals 65 14.2 3416 706 459 2420
Body 92 13.12 4988 897 679 3462
Food 53 14.25 3046 519 379 2066
Cloth 44 12.95 2296 450 273 1631
House 31 13.81 1784 267 234 1205
Agriculture 36 13.81 2285 409 270 1337
Action 39 14.08 2336 362 353 1436
Motion 47 15.36 2910 465 488 1617
Possession 33 12.97 1841 237 313 1175
Space 59 12.8 3544 406 529 1970
Quantity 29 12.69 1456 166 296 924
Time 42 12.05 2711 289 370 1455
Sense 34 15.38 1975 292 426 1337
Values 42 13.17 2243 346 319 1670
Cognit 38 13.45 2299 277 323 1332
Language 31 14.13 1978 277 271 1073
Society 27 14.19 1353 229 183 957
Warfare 26 14.46 1428 277 158 940
Law 16 15.19 1078 135 135 682
Religion 11 14.36 592 104 72 440
Overall 881 13.94 51700 8097 7309 32406
Table 5: The data overview(frequency) of our dataset.
Appendix B Experiment Settings
B.1 Template Using
The templates used in this work are shown in Table 6. Notably, the CN-t condition also employs English prompts.
Language Mode Template
EN,CN-t Base When "{cue_word}" is mentioned, people often think of the following words:
CSP You are a person with {culture} cultural background. You will be performing a word association task. Please directly answer the association word. When "{cue_word}" is mentioned, people often think of the following words:
CCT You are a person with {culture} cultural background. You will be performing a word association task. Please directly answer the association word. Before you respond, think about how {culture} culture is different from {cultures} cultures. When "{cue_word}" is mentioned, people often think of the following words
CN Base {CJK*}UTF8gbsn当提起"{cue_word}",人们往往会想到的词是:
CSP {CJK*}UTF8gbsn你是一个中国文化背景的人。 {CJK*}UTF8gbsn你将进行词联想任务,请直接说出你联想到的词。 {CJK*}UTF8gbsn当提起"{cue_word}",人们往往会想到的词是:
CCT {CJK*}UTF8gbsn你是一个中国文化背景的人。 {CJK*}UTF8gbsn你将进行词联想任务, 请直接说出你联想到的词。 {CJK*}UTF8gbsn在回答之前, 请你注意中国文化与美国、英国、大洋洲文化的不同。 {CJK*}UTF8gbsn当提起"{cue_word}",人们往往会想到的词是:
Table 6: Templates for different tasks and regions, where red denotes additional cultural information, blue indicates cross-cultural thinking instructions, and orange marks task descriptions aimed at preventing LLMs from generating redundant responses.
B.2 Hyperparaments
The detailed setting of hyperparaments are shown in Table  7
Phase Hyperparameters Value
Train learn rate 1e-4
epsilon 1e-3
max length 64
l2 rate 0
LoRA rank 8
LoRA alpha 16
LoRA dropout 0.05
batch size 8
epoch 5
Generate max new tokens 5
temperature 1
epsilon 1e-3
Table 7: Relative hyperparameters
Appendix C Language Ablation Study
Llama(baseline)
PWR@K 3 5 10 20
CN 7.31 11.03 18.43 24.01
CN-t 5.51 8.79 15.62 23.75
Qwen(baseline)
CN 8.61 11.65 17.16 24.11
CN-t 5.55 7.21 10.33 15.26
Table 8: Comparison results after translating Chinese prompts and lexical items into English. Bold values indicate stronger activation of word association capabilities in Chinese linguistic contexts.
In § 5, we performed an ablation study introducing a CN-t (Chinese-to-English translated) group to investigate the impact of linguistic differences on our results, controlling for language while preserving English context. Chinese-associated words were translated using MUSE bilingual dictionaries Lample et al. (2017), with Opus-MT-zh-en Tiedemann et al. (2024) covering uncovered pairs. Results in Table 8 suggest that Chinese contexts more effectively activate Chinese word association capabilities, indicating that the main experimental performance differences stem from cultural rather than linguistic factors.
Appendix D Granular Performance
The granular performance is shown in Figure 5. When K=3, 5, 10, the CultureSteer results consistently exhibit comparably weaker performance in Global Knowledge categories, while cultural disparities in Llama remain greater than those in Qwen.
(a) Performance in PWR@3
(b) Performance in PWR@5
(c) Performance in PWR@10
Figure 5: Graunalr performance in PWR@3, 5 and 10